# Reinforcement Learning with Action Sequence for Data-Efficient Robot Learning

### v1

@@ 3.2 Action Execution and Training Details
While the idea of using action sequence is simple, there are two important yet small details for effectively training RL agents with action sequence: (i) how we execute actions at each time step to control robots and (ii) how we store training data and sample batches for training.
@@
With the policy that outputs an action sequence $\mathbf{a}_{t:t+K}$, one important question is how to execute actions at time step $i\in\{t,...,t+K-1\}$. For this, we use the idea of Zhao et al. (2023) that utilizes temporal ensemble, which computes $\mathbf{a}_{t:t+K}$ every time step, saves it to a buffer, and executes a weighted average $\sum_{i}w_{i}\mathbf{a}_{t-i}/\sum w_{i}$ where $w_{i}=\exp(-m*i)$ denotes a weight that assigns higher value to more recent actions. We find this scheme outperforms the alternative of computing $\mathbf{a}_{t:t+K}$ every $K$ steps and executing each action for subsequent $K$ steps on most tasks we considered, except on several tasks that need reactive control.
Storing training data from environment interaction
When storing samples from online environment interaction, we store a transition $(\mathbf{o}_{t},\hat{\mathbf{a}}_{t},r_{t+1},\mathbf{o}_{t+1})$ where $\hat{\mathbf{a}}_{t}$ denotes an action executed at time step $t$. For instance, if we use temporal ensemble for action execution, $\hat{\mathbf{a}}_{t}$ is a weighted average of action outputs obtained from previous $K$ time steps.
@@
When sampling training data from the replay buffer, we sample a transition with action sequence, i.e., $(\mathbf{o}_{t},\hat{\mathbf{a}}_{t:t+K},r_{t+1},\mathbf{o}_{t+1})$. If we sample time step $t$ near the end of episode so that we do not have enough data to construct a full action sequence, we fill the action sequence with null actions. In particular, in position control where we specify the position of joints or end effectors, we repeat the action from the last step so that the agent learns not to change the position. On the other hand, in torque control where we specify the force to apply to joints, we set the action after the last step to zero so that agent learns to not to apply force.
Figure 3: Examples of robotic tasks. We study CQN-AS on 53 robotic tasks spanning mobile bi-manual manipulation, whole-body control, and tabletop manipulation tasks from BiGym (Chernyadev et al., 2024), HumanoidBench (Sferrazza et al., 2024), and RLBench (James et al., 2020).
@@ 4 Experiment
    What is the contribution of each component in CQN-AS?
@@
    Under which conditions is CQN-AS effective? When does CQN-AS fail?
Baselines for fine-grained control tasks with demonstrations
For tasks that need high-precision control, e.g., manipulation tasks from BiGym and RLBench, we consider model-free RL baselines that aim to learn deterministic policies, as we find that stochastic policies struggle to solve such fine-grained control tasks. Specifically, we consider (i) Coarse-to-fine Q-Network (CQN; Seo et al. 2024), a value-based RL algorithm that learns to zoom-into continuous action space in a coarse-to-fine manner, and (ii) DrQ-v2+, an optimized demo-driven variant of a model-free actor-critic algorithm DrQ-v2 (Yarats et al., 2022) that uses a deterministic policy algorithm and data augmentation. We further consider (iii) Action Chunking Transformer (ACT; Zhao et al. 2023), a BC algorithm that trains a transformer (Vaswani et al., 2017) policy to predict action sequence and utilizes temporal ensemble for executing actions, as our highly-optimized BC baseline.
Baselines for whole-body control tasks with dense reward
For locomotion tasks with dense reward, we consider (i) Soft Actor-Critic (SAC; Haarnoja et al. 2018), a model-free actor-critic RL algorithm that maximizes action entropy, and (ii) Coarse-to-fine Q-Network (CQN; Seo et al. 2024). Moreover, although it is not the goal of this paper to compare against model-based RL algorithms, we also consider two model-based baselines: (iii) DreamerV3 (Hafner et al., 2023), a model-based RL algorithm that learns a latent dynamics model and a policy using imagined trajectories and (iv) TD-MPC2 (Hansen et al., 2024), a model-based RL algorithm that learns a latent dynamics model and utilizes local trajectory optimization in imagined latent trajectories.
@@
For training with expert demonstrations, we follow the setup of Seo et al. (2024). Specifically, we keep a separate replay buffer that stores demonstrations and sample half of training data from demonstrations. We also relabel successful online episodes as demonstrations and store them in the demonstration replay buffer. For CQN-AS, we use an auxiliary BC loss from Seo et al. (2024) based on large margin loss (Hester et al., 2018). For actor-critic baselines, we use an auxiliary BC loss that minimizes L2 loss between the policy outputs and expert actions. We implement CQN-AS based on a publicly available source code of CQN [^2] https://github.com/younggyoseo/CQN based on PyTorch (Paszke et al., 2019). We will release the full source code upon publication.
Figure 4: BiGym results on 25 sparsely-rewarded mobile bi-manual manipulation tasks (Chernyadev et al., 2024). All experiments are initialized with 17 to 60 human-collected demonstrations, and RL methods are trained with an auxiliary BC objective. On many of the challenging long-horizon tasks, CQN-AS quickly learns to match the performance of ACT (Zhao et al., 2023) and surpass it through online learning, while other RL baselines fail to effectively accelerate training with noisy human-collected demonstrations. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
Figure 5: HumanoidBench results on 8 densely-rewarded humanoid control tasks (Sferrazza et al., 2024). All the experiments start from scratch and RL methods do not have an auxiliary BC objective. CQN-AS significantly outperforms other model-free RL baselines on most tasks. CQN-AS often achieves competitive performance to model-based RL baselines, which is intriguing but not the main goal of this paper. For CQN-AS and CQN, we report the results aggregated over 4 runs. For other baselines, we report the results aggregated over 3 runs available from public website. The solid line and shaded regions represent the mean and confidence intervals.
@@ 4.1 BiGym Experiments
We study CQN-AS on mobile bi-manual manipulation tasks from BiGym (Chernyadev et al., 2024). BiGym’s human-collected demonstrations are often noisy and multi-modal, posing challenges to RL algorithms which should effectively leverage demonstrations for solving sparsely-rewarded tasks.
@@
Because we find that not all demonstrations from BiGym benchmark can be successfully replayed [^3] We use demonstrations available at the date of Oct 1st with the commit 018f8b2., we replay all the demonstrations and only use the successful ones as demonstrations. We do not discard ones that fail to be replayed, but we use them as training data with zero reward because they can still be useful as failure experiences. To avoid training with too few demonstrations, we exclude the tasks where the ratio of successful demonstrations is below 50%. This leaves us with 25 tasks, each with 17 to 60 demonstrations. For visual observations, we use RGB observations with 84$\times$84 resolution from head, left_wrist, and right_wrist cameras. We also use low-dimensional proprioceptive states from proprioception, proprioception_grippers, and proprioception_floating_base sensors. We use (i) absolute joint position control action mode and (ii) floating base that replaces locomotion with classic controllers. We use the same set of hyperparameters for all the tasks, in particular, we use action sequence of length 16. More details on BiGym experiments are available in Appendix A.
@@
Figure 4 shows the experimental results on BiGym benchmark. We find that CQN-AS quickly matches the performance of ACT and outperforms it through online learning on 20/25 tasks, while other RL algorithms fail to do so especially on challenging long-horizon tasks such as Move Plate and Saucepan To Hob. A notable result here is that CQN-AS enables solving these challenging BiGym tasks while other RL baselines, i.e., CQN and DrQ-v2+, completely fail as they achieve 0% success rate. This result highlights the capability of CQN-AS to accelerate RL training from noisy, multi-modal demonstrations collected by humans.
@@
However, we find that CQN-AS struggles to achieve meaningful success rate on some of the long-horizon tasks that require interaction with delicate objects such as cup or cutlery, leaving room for future work to incorporate advanced vision encoders (He et al., 2016; 2022) or critic architectures (Kapturowski et al., 2023; Chebotar et al., 2023; Springenberg et al., 2024).
@@ 4.2 HumanoidBench Experiments
To show that CQN-AS can be generally applicable to tasks without demonstrations, we study CQN-AS on densely-rewarded humanoid control tasks from HumanoidBench (Sferrazza et al., 2024).
@@
For HumanoidBench, we follow a standard setup that trains RL agents from scratch, which is also used in original benchmark (Sferrazza et al., 2024). Specifically, we use low-dimensional states consisting of proprioception and privileged task information as inputs. For tasks, we simply select the first 8 locomotion tasks in the benchmark. Following the original benchmark that trains RL agents for environment steps that roughly requires 48 hours of training, we report the results of CQN-AS and CQN for 7 million steps. For baselines, we use the results available from the public repository, which are evaluated on tasks with dexterous hands, and we also evaluate our algorithm on tasks with hands. We use the same set of hyperparameters for all the tasks, in particular, we use action sequence of length 4. More details on HumanoidBench experiments are available in Appendix A.
Comparison to model-free RL baselines
Figure 5 shows the results on on HumanoidBench. We find that, by learning the critic network with action sequence, CQN-AS outperforms other model-free RL baselines, i.e., CQN and SAC, on most tasks. In particular, the difference between CQN-AS and baselines becomes larger as the task gets more difficult, e.g., baselines fail to achieve high episode return on Walk and Run tasks but CQN-AS achieves strong performance. This result shows that our idea of using action sequence can be applicable to generic setup without demonstrations.
CQN-AS often achieves competitive performance to model-based RL baselines
While outperforming model-based RL algorithms is not the goal of this paper, we find that CQN-AS often achieves competitive performance to model-based RL baselines, i.e., DreamerV3 and TD-MPC2, on tasks such as Run or Sit Simple. This result shows the potential of our idea to enable RL agents to learn useful value functions on challenging tasks, without the need to explicitly learn dynamics model. We also note that incorporating our idea into world model learning could be an interesting direction.
@@ 4.3 RLBench Experiments
For RLBench experiments, we use the official CQN implementation for collecting demonstrations and reproducing the baseline results on the same set of tasks. Specifically, we use RGB observations with 84$\times$84 resolution from front, wrist, left_shoulder, and right_shoulder cameras. We also use low-dimensional proprioceptive states consisting of 7-dimensional joint positions and a binary value for gripper open. We use 100 demonstrations and delta joint position control action mode. We use the same set of hyperparameters for all the tasks, in particular, we use action sequence of length 4. More details on RLBench experiments are available in Appendix A.
@@
Because RLBench provides synthetic clean demonstrations, as we expected, Figure 6 shows that CQN-AS achieves similar performance to CQN on many of the tasks, except 2/25 tasks where it hurts the performance. But we still find that CQN-AS achieves quite superior performance to CQN on some challenging long-horizon tasks such as Open Oven or Take Plate Off Colored Dish Rack. These results, along with results from BiGym and HumanoidBench, show that CQN-AS can be used in various benchmark with different characteristics.
@@ 4.4 Ablation Studies, Analysis, Failure Cases
Figure 7(a) shows the performance of CQN-AS with different action sequence lengths on two BiGym tasks. We find that training the critic network with longer action sequences improves and stabilizes performance.
@@
Figure 7(b) shows the performance of CQN-AS without RL objective that trains the model only with BC objective on successful demonstrations. We find this baseline significantly underperforms CQN-AS, which shows that RL objective is indeed enabling the agent to learn from online trial-and-error experiences.
@@
Figure 7(c) shows that performance largely degrades without temporal ensemble on Saucepan To Hop as temporal ensemble induces a smooth motion and thus improves performance in fine-grained control tasks. But we also find that temporal ensemble can be harmful on Reach Target Single. We hypothesize this is because temporal ensemble often makes it difficult to refine behaviors based on recent visual observations. Nonetheless, we use temporal ensemble for all the tasks as we find it helps on most tasks and we aim to use the same set of hyperparameters.
@@
Figure 7(d) shows that CQN-AS underperforms CQN on locomotion tasks with torque control. We hypothesize this is because a sequence of joint positions usually has a semantic meaning in joint spaces, making it easier to learn with, when compared to learning how to apply a sequence of torques. Addressing this failure case is an interesting future direction.
(a) Effect of action sequence length
(b) Effect of RL objective
(c) Effect of temporal ensemble
(d) Failure mode: Torque control
Figure 7: Ablation studies and analysis on the effect of (a) action sequence, (b) RL objective, and (c) temporal ensemble. (d) We also provide results on locomotion tasks from DeepMind Control Suite (Tassa et al., 2020), where CQN-AS fails to improve performance. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
@@ 5 Related Work
Recent behavior cloning approaches have shown that predicting a sequence of actions enables the policy to effectively imitate noisy expert trajectories and helps in dealing with idle actions from human pauses during data collection (Zhao et al., 2023; Chi et al., 2023). In particular, Zhao et al. (2023) train a transformer model (Vaswani et al., 2017) that predicts action sequence and Chi et al. (2023) train a denoising diffusion model (Ho et al., 2020) that approximates the action distributions. This idea has been extended to multi-task setup (Bharadhwaj et al., 2024), mobile manipulation (Fu et al., 2024b) and humanoid control (Fu et al., 2024a). Our work is inspired by this line of work and proposed to learn RL agents with action sequence.
@@
In the context of reinforcement learning, Medini & Shrivastava (2019) proposed to pre-compute frequent action sequences from expert demonstrations and augment the action space with these sequences. However, this idea introduces additional complexity and is not scalable to setups without demonstrations. One recent work relevant to ours is Saanum et al. (2024) that encourages a sequence of actions from RL agents to be predictable and smooth. But this differs from our work in that it uses the concept of action sequence only for computing the penalty term. Recently, Ankile et al. (2024) point out that RL with action sequence is challenging and instead proposes to use RL for learning a single-step policy that corrects action sequence predictions from BC. In contrast, our work shows that training RL agents with action sequence is feasible and leads to improved performance compared to prior RL algorithms.

### v2

@@ 3.2 Action Execution and Training Details
With the policy that outputs an action sequence $a_{t:t+K}$, one question is how to execute actions at time step $i\in\{t,...,t+K-1\}$. For this, we use temporal ensemble (Zhao et al., 2023) that computes $a_{t:t+K}$ every time step, saves it to a buffer, and executes a weighted average $\sum_{i}w_{i}\bar{a}_{t}^{i}/\sum w_{i}$ where $\bar{a}_{t}^{i}$ denotes an action for step $t$ computed at step $t-i$, $w_{i}=\exp(-m*i)$ denotes a weight that assigns higher value to more recent actions, with $m$ as a hyperparameter that adjusts the weighting magnitude. We find this scheme outperforms the alternative of computing $a_{t:t+K}$ every $K$ steps and executing each action for subsequent $K$ steps on most tasks we considered, except on several tasks that need reactive control.
Storing training data
When storing samples from the environment, we store a transition $(\mathbf{o}_{t},\hat{a}_{t},r_{t+1},\mathbf{o}_{t+1})$ where $\hat{a}_{t}$ denotes an action executed at time step $t$. For instance, if we use temporal ensemble for action execution, $\hat{a}_{t}$ is a weighted average of action outputs obtained from previous $K$ time steps, i.e., $\hat{a}_{t}=\sum_{i}w_{i}\bar{a}_{t}^{i}/\sum w_{i}$.
@@
When sampling training data from the replay buffer, we sample a transition with action sequence, i.e., $(\mathbf{o}_{t},\hat{a}_{t:t+K},r_{t+1},\mathbf{o}_{t+1})$. If we sample time step $t$ near the end of episode so that we do not have enough data to construct a full action sequence, we fill the action sequence with null actions. In particular, in position control where we specify the position of joints or end effectors, we repeat the action from the last step so that the agent learns not to change the position. In torque control where we specify the force to apply, we set the action after the last step to zero so that agent learns to not to apply force.
Figure 5: BiGym results on 25 sparsely-rewarded mobile bi-manual manipulation tasks. All experiments are initialized with 17 to 60 human-collected demonstrations, and RL methods are trained with an auxiliary BC objective. On many tasks, CQN-AS quickly matches the performance of ACT (Zhao et al., 2023) and surpasses it through online learning, also outperforming other RL baselines. We report the success rate over 25 episodes. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 8 runs.
@@ 4 Experiment
    What is the effect of each component in CQN-AS?
@@
    Under which conditions is CQN-AS effective?
Baselines for tasks with demonstrations
For manipulation tasks from BiGym and RLBench, we consider model-free RL baselines that learn deterministic policies, as we find that stochastic policies struggle to solve such fine-grained control tasks. Specifically, we consider (i) Coarse-to-fine Q-Network (CQN; Seo et al. 2024), our backbone algorithm and (ii) DrQ-v2+, an optimized demo-driven variant of an actor-critic algorithm DrQ-v2 (Yarats et al., 2022) that uses a deterministic policy algorithm and data augmentation. We further consider (iii) Action Chunking Transformer (ACT; Zhao et al. 2023), a BC algorithm that trains a transformer (Vaswani et al., 2017) policy to predict action sequence and utilizes temporal ensemble, as our BC baseline.
Baselines for humanoid control tasks with dense reward
For HumanoidBench tasks, we consider (i) Soft Actor-Critic (SAC; Haarnoja et al. 2018), a model-free actor-critic RL algorithm that maximizes action entropy and (ii) CQN.
@@
For training with expert demonstrations, we follow the setup of Seo et al. (2024). We keep a separate replay buffer that stores demonstrations and sample half of training data from demonstrations. We also relabel successful online episodes as demonstrations and store them in the demonstration replay buffer. For CQN-AS, we use an auxiliary BC loss from Seo et al. (2024) based on large margin loss (Hester et al., 2018). For actor-critic baselines, we use an auxiliary BC loss that minimizes L2 loss between the policy outputs and expert actions.
Figure 6: HumanoidBench results on 8 densely-rewarded humanoid control tasks (Sferrazza et al., 2024). All the experiments start from scratch and RL methods do not have an auxiliary BC objective. CQN-AS significantly outperforms other model-free RL baselines on most tasks. For CQN-AS and CQN, we report the results aggregated over 8 runs. For SAC, we report the results aggregated over 3 runs available from public website. The solid line and shaded regions represent the mean and confidence intervals.
@@ 4.1 BiGym Experiments
We study CQN-AS on mobile bi-manual manipulation tasks from BiGym (Chernyadev et al., 2024). BiGym’s human-collected demonstrations are often noisy and multi-modal, posing challenges to RL algorithms which should leverage demonstrations for solving sparsely-rewarded tasks.
@@
We consider 25 BiGym tasks with 17 to 60 demonstrations. We use RGB observations with 84$\times$84 resolution from head, left_wrist, and right_wrist cameras. We also use low-dimensional proprioceptive states. We use (i) absolute joint position control action mode and (ii) floating base that replaces locomotion with classic controllers. We use the same set of hyperparameters for all the tasks. Details on BiGym experiments are available in Appendix A.
@@
Figure 5 shows that CQN-AS quickly matches the performance of ACT and outperforms it through online learning on most tasks, while other RL algorithms fail to do so especially on challenging long-horizon tasks such as Move Plate and Saucepan To Hob. A notable result here is that CQN-AS enables solving challenging BiGym tasks while other RL baselines completely fail as they achieve 0% success rate on many tasks.
@@
However, CQN-AS struggles to achieve meaningful success rate on some of the long-horizon tasks that require interaction with delicate objects such as cup or cutlery. This leaves room for future work to incorporate advanced vision encoders (He et al., 2016) or critic architectures (Chebotar et al., 2023; Springenberg et al., 2024).
@@ 4.2 HumanoidBench Experiments
To show that CQN-AS is generally applicable to tasks without demonstrations, we study CQN-AS on densely-rewarded tasks from HumanoidBench (Sferrazza et al., 2024).
@@
We follow a standard setup that trains RL agents from scratch. We use low-dimensional states consisting of proprioception and privileged task information as inputs. For tasks, we simply select the first 8 locomotion tasks in the benchmark. For baselines, we use the results available from the public repository, which are evaluated on tasks with dexterous hands, and we also evaluate our algorithm on tasks with hands. We use the same set of hyperparameters for all the tasks. More details are available in Appendix A.
Comparison to baselines
Figure 6 shows that, by learning the critic network with action sequence, CQN-AS outperforms other model-free RL baselines, i.e., CQN and SAC, on most tasks. In particular, the difference between CQN-AS and baselines becomes larger as the task gets more difficult, e.g., baselines fail to achieve high episode return on Walk and Run tasks but CQN-AS achieves strong performance. This result shows that our idea of using action sequence can be applicable to generic setup without demonstrations.
@@ 4.3 RLBench Experiments
We use the official CQN implementation for collecting demonstrations and reproducing the baseline results on the same set of tasks. We use RGB observations with 84$\times$84 resolution from front, wrist, left_shoulder, and right_shoulder cameras. We also use low-dimensional proprioceptive states consisting of 7-dimensional joint positions and a binary value for gripper open. We use 100 demonstrations and delta joint position control action mode. We use the same set of hyperparameters for all the tasks, in particular, we use action sequence of length 4. More details on RLBench experiments are available in Appendix A.
@@
Because RLBench provides synthetic clean demonstrations, as we expected, Figure 7 shows that CQN-AS achieves similar performance to CQN on most tasks, except 2/25 tasks where it hurts the performance. But we still find that CQN-AS achieves quite superior performance to CQN on some challenging long-horizon tasks such as Open Oven or Take Plate Off Colored Dish Rack. These results show that CQN-AS can be used in various benchmark with different characteristics.
(a) Effect of action sequence length
(b) Effect of RL objective
(c) Effect of $N$-step
(d) Effect of temporal ensemble
(e) Effect of temporal ensemble magnitude $m$
(f) Failure mode: Torque control
Figure 8: Ablation studies and analysis on the effect of (a) action sequence, (b) RL objective, (c) $N$-step return, and (d & e) temporal ensemble. (f) We also provide results on locomotion tasks from DeepMind Control Suite (Tassa et al., 2020), where CQN-AS fails to improve performance. The solid line and shaded regions represent the mean and confidence intervals, respectively, across 4 runs.
@@ 4.4 Ablation Studies, Analysis, Failure Cases
Figure 8(a) shows the performance of CQN-AS with different action sequence lengths. We find that training the critic network with longer action sequences improves performance.
@@
Figure 8(b) shows the performance of CQN-AS without RL objective that trains the model only with BC objective on successful demonstrations. We find this baseline significantly underperforms CQN-AS, which shows that RL objective enables the agent to learn from trial-and-error experiences.
Effect of $N$-step return
Figure 8(c) shows experimental results with varying $N$-step returns. We find that too high $N$-step return significantly degrades performance. We hypothesize this is because the variance from $N$-step return makes it difficult to learn useful value functions.
@@
Figure 8(d) shows that performance degrades without temporal ensemble on Saucepan To Hop as temporal ensemble induces a smooth motion and thus improves performance in fine-grained control tasks. But we also find that temporal ensemble can be harmful on Reach Target Single. We hypothesize this is because temporal ensemble uses predictions from previous steps and thus makes it difficult to refine behaviors based on recent visual observations. Nonetheless, we use temporal ensemble for all the tasks as we find it helps on most tasks and we aim to use the same set of hyperparameters.
Effect of temporal ensemble magnitude
We further provide results with different temporal ensemble magnitudes by adjusting a hyperparameter $m$ in Figure 8(e). Here, higher $m$ puts higher weights on recent actions and thus very high $m$ corresponds to using only first action. Similarly to previous paragraph, we find that higher $m$ leads to better performance on Reach Target Single that needs fast reaction, but degrades performance on Saucepan To Hop.
@@
Figure 8(f) shows that CQN-AS underperforms CQN on locomotion tasks with torque control. We hypothesize this is because a sequence of joint positions usually has a semantic meaning in joint spaces, making it easier to learn with, when compared to learning how to apply a sequence of torques. Addressing this failure case is an interesting future direction.
@@ 5 Related Work
Recent behavior cloning approaches have shown that predicting a sequence of actions enables the policy to imitate noisy expert trajectories and helps in dealing with idle actions from human pauses during data collection (Zhao et al., 2023; Chi et al., 2023). Zhao et al. (2023) train a transformer model (Vaswani et al., 2017) that predicts action sequence and Chi et al. (2023) train a denoising diffusion model (Ho et al., 2020) that approximates the action distributions. This idea has been extended to multi-task setup (Bharadhwaj et al., 2024), mobile manipulation (Fu et al., 2024b) and humanoid control (Fu et al., 2024a). Our work is inspired by this line of work and proposed to learn RL agents with action sequence.
@@
In the context of RL, Medini & Shrivastava (2019) propose to pre-compute frequent action sequences from expert demonstrations and augment the action space with these sequences. However, this idea introduces additional complexity and is not scalable to setups without demonstrations. One recent work relevant to ours is Saanum et al. (2024) that encourage a sequence of actions from RL agents to be predictable and smooth. But this differs from our work in that it uses action sequence only for computing the penalty term. Recently, Ankile et al. (2024) point out that RL with action sequence is challenging and instead propose to use RL for learning a single-step policy that corrects action sequence predictions from BC. In contrast, we show that RL with action sequence is feasible and improves performance of RL algorithms.